Architecture. Architects often use the equation x + y = 18 to determine the height y, in inches, of the riser of a step when the tread is x inches wide. Express the slope of stairs designed with this equation without using the variable y.
Answer by richard1234(7193)

If the height is the dependent variable, then

y = 18-x = -x+18

Slope is -1.